OpenStudy (anonymous):

6[-6x=6+4(x+1)]=6x+2

OpenStudy (anonymous):

-6x=6+4x+4=6x+2/6 -6x=6+4x+4,-6x=6x+2/6,6+4x+4=6x+2/6 like this